Dr Paul Coddington Northeast Parallel Architectures Center 111 College Place Syracuse University Syracuse, NY 13244 U.S.A.

Phone: (315) 443-4883 FAX: (315) 443-1973 Email: paulc@npac.syr.edu

April 19, 1994

To whom it may concern,

Kamala Anupindi is a graduate student working for me at the Northeast Parallel Architectures Center at Syracuse University. I am hoping that she will be able to attend a conference in Europe in the summer (CONPAR 94 in Austria), however in order to do that she needs an extension of her visa to allow her to return to the United States to continue her graduate studies. I hope that this can be achieved as soon as possible.

Yours sincerely,